Let the coin is tossed n times.
∵p (at least 2 heads) =1−[p (one heads) +p (No heads) }
As we know, by binominal probability theorem the probability of getting r success in n trials with p being the probability of success and q be the probability of failure, is given by nCr(p)r(q)n−r .
Let head be considered as the success and tail be the failure probability of getting head in a toss =p=21 and probability of getting tail in a toss =q=21
amp;∴P( one head )=nC1(21)1(21)n−1amp;=nC1(21)namp;⇒P (one head) =n(21)namp; Similarly, P(No heads)=nC0(21)0(21)namp;⇒P( No heads )=(21)n{∵nco=1}amp;∴P (at least 2 heads) =1−{n(21)n+(21)n}
amp;=1−2n(n+1)amp;∵P( at least 2 heads )≥0.96amp;⇒1−2n(n+1)≥0.96amp;⇒0.04≥2nn+1amp;⇒251≥2nn+1amp;⇒n+12n≥25amp;⇒n≥8
So, the minimum number of times a fair coin to be tossed is 8.
Hint:
(i) P (at least 2 heads) =1−{P (one head} + P (No heads)}
(ii) Use binomial probability theorem and simplify it.